OpenStudy (anonymous):

What is the result of isolating y2 in the equation below? (x - 2)2 + y2 = 64 A. y2 = -x2 + 4x + 4 B. y2 = 64 - x2 C. y2 = x2 + 4x + 60 D. y2 = -x2 + 4x + 60

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

\[\left( a-b \right)^2=a^2+b^2-2ab\]

OpenStudy (anonymous):

wut??

Directrix (directrix):

(x - 2)^2 + y^2 = 64 Subtract (x - 2)^2 from both sides of the equation. (x - 2)^2 + y^2 = 64 - (x - 2)^2 -(x - 2)^2 ------------------------------

Directrix (directrix):

y^2 = 64 - (x - 2)^2 Next Step @2kende What does this equal when you simplify? --> 64 - (x - 2)^2 =